A farmer plans to cover a 70 yd by 125 yd rectangular field with enriched soil 3 inches deep. Each truckload of soil costs $280 and has volume 85 cubic yards. How many truckloads of soil does he need? What is the total cost of the soil?
Answer by rfer(16322)

.25/3=0.083
70*125*0.083=726.25 cu tds
726.25/85=8.54 truckloads
8.54*280=2391.20